# Input Convex Lipschitz RNN: A Fast and Robust Approach for Engineering Tasks

### v3

@@ Abstract
Computational efficiency and non-adversarial robustness are critical factors in real-world engineering applications. Yet, conventional neural networks often fall short in addressing both simultaneously, or even separately. Drawing insights from natural physical systems and existing literature, it is known that an input convex architecture enhances computational efficiency, while a Lipschitz-constrained architecture bolsters non-adversarial robustness. By leveraging the strengths of convexity and Lipschitz continuity, we develop a novel network architecture, termed Input Convex Lipschitz Recurrent Neural Networks. This model is explicitly designed for fast and robust optimization-based tasks and outperforms existing recurrent units across a spectrum of engineering tasks in terms of computational efficiency and non-adversarial robustness, including real-world solar irradiance prediction for Solar PV system planning at LHT Holdings in Singapore and real-time Model Predictive Control optimization for a nonlinear chemical reactor.
Index Terms:
Deep Learning, Lipschitz Constrained Neural Networks, Input Convex Neural Networks, Computational Efficiency, Non-Adversarial Robustness, Model Predictive Control, Solar Irradiance Forecasting
I Introduction
Two crucial criteria for the resolution of real-world engineering challenges are computational efficiency and non-adversarial robustness. This big-data era has empowered the capabilities of neural networks and transformed them into an incredibly powerful tool, offering comprehensive solutions to numerous engineering problems such as image classification and process modeling. This work draws inspiration from nature-inspired design to develop a fast and robust solution to model nonlinear systems in various engineering disciplines.
Model Predictive Control (MPC) represents an advanced control technique applied to manage a process while complying with a set of constraints, akin to an optimization challenge. Traditional MPC based on first-principles models encounters limitations, particularly in scenarios with intricate dynamics, where deriving these models proves infeasible. Consequently, researchers proposed neural network-based MPC (NN-MPC) as a viable alternative to model system dynamics [1, 2, 3, 4, 5, 6, 7, 8]. However, MPC built on conventional neural networks often falls short in meeting the aforementioned critical criteria. Nonetheless, efforts have been made to mitigate these limitations. For example, the computational efficiency of NN-MPC is enhanced by implementing an input convex architecture in neural networks [9]. Additionally, non-adversarial robustness of NN-MPC is improved by developing Lipschitz-constrained neural networks (LNNs) [10].
In real-world applications, computational efficiency plays a pivotal role due to the imperative need for real-time decision-making, especially in critical engineering processes such as chemical process operations [9]. We define computational efficiency as the speed at which computational tasks are completed within a given execution timeframe. An effective approach to achieve this efficiency is through the transformation from non-convex to convex structures. Convexity is an ubiquitous characteristic observed in various physical systems. In energy potentials, the potential energy function of certain systems, such as simple harmonic oscillators, can exhibit a convex behavior. For example, the potential energy function of a mass on a spring is a classic example of a convex function in physics [11, 12]. In the context of magnetic fields, the region around a single magnetic pole, where another magnetic pole encounters an attractive force, is considered convex [13, 14]. In the theory of general relativity, the curvature of space-time around massive objects like stars or black holes can be described as convex in certain regions. This curvature influences the paths of objects and light rays near these massive bodies [15].
@@ I Introduction
Building upon nature-inspired design and extending its application to neural networks, Input Convex Neural Networks (ICNNs) is a class of neural network architectures intentionally crafted to maintain convexity in their output with respect to the input. Leveraging these models could be particularly advantageous in various engineering problems, notably in applications such as MPC, due to their inherent benefits especially in optimization. This concept was first introduced by [16] as Feedforward Neural Networks (FNN), and subsequently extended to Recurrent Neural Networks (RNN) by [17], and Long Short-Term Memory (LSTM) by [9].
@@
Additionally, in real-world engineering applications, non-adversarial robustness plays a critical role due to the prevalent noise present in most real-world data, which significantly hampers neural network performance. Our definition of robustness aligns closely with the principles outlined by [10]. Given the inherent noise in real-world process data, our goal is to improve neural networks by effectively learning from noisy training data for comprehensive end-to-end applications. Leveraging LNNs offers a potential solution to the robustness challenge. This concept was first introduced by [18] as FNN, and subsequently extended to RNN by [19], and Convolutional Neural Networks (CNN) by [20].
@@
Motivated by the above considerations, in this work, we introduce an Input Convex Lipschitz RNN (ICLRNN) that amalgamates the strengths of ICNNs and LNNs. This approach aims to achieve the concurrent resolution of computational efficiency and non-adversarial robustness, which remains a challenge in real-world engineering applications. The core motivation is to develop a fast and robust neural network tailored for optimization frameworks, as shown in Fig. 1. In this study, we validate the efficacy of ICLRNN across various engineering problems that include time series forecasting and neural network-based MPC. Specifically, for time-series forecasting, we engage in planning for Solar Photovoltaic (PV) systems with real-time data from LHT Holdings plants in Singapore, which can be further leveraged for energy dispatch optimization. Furthermore, we use a nonlinear chemical reactor example to show the benefits of ICLRNN in the context of MPC.
Figure 1: System architecture.
@@
In summary, Section II offers an overview of modern high-performing recurrent units, ICNNs and LNNs. Section III presents the proposed ICLRNN, substantiating its theoretical attributes as input convex and Lipschitz continuous. Section IV illustrates that ICLRNN surpasses state-of-the-art recurrent units in various engineering tasks, including real-time solar irradiance forecasting on Solar PV systems, and optimization and control of chemical processes using MPC. Section V delves into the potential limitations of ICLRNN uncovered through experiments conducted on the Long Range Arena benchmark [21].
II Background
II-A Notations
@@ II-B High-Performing RNN
In recent years, several high-performing recurrent units have emerged, including UniCORNN [22], rank-coding based RNNs [23], and Linear Recurrent Units [24]. However, these variants primarily emphasize enhancing accuracy, especially for very long time-dependent sequences modelling, which often at the expense of model simplicity and computational efficiency. However, in many engineering applications that we are currently working on with industry partners, accuracy alone does not dictate the suitability of a model. In fact, from an engineering point of view, simplicity and robustness are also important in addition to accuracy, as they ensure the model can be easily adapted and incorporated into the existing industrial operational systems, and can be trusted in an uncertain environment with noise and disturbances. Therefore, our goal is to obtain a model that strikes a balance between these factors while maintaining a satisfactory level of accuracy. This motivation underpins our investigation into the fusion of ICNNs and LNNs.
II-C Input Convex RNN
@@
The initiative behind ICNNs is to leverage the power of neural networks, specifically designed to maintain convexity within their decision boundaries [16]. ICNNs aim to combine the strengths of neural networks in modeling complex data with the advantages of convex optimization, which ensures the convergence to a global optimal solution. They are particularly attractive for control and optimization problems, where achieving globally optimal solutions is essential. By maintaining convexity, ICNNs help ensure that the optimization problems associated with neural networks (e.g., neural network-based optimization) remain tractable, thus addressing some of the challenges of non-convexity and the potential for suboptimal local solutions in traditional neural networks.
A foundational work, known as Input Convex RNN (ICRNN), serves as one of the baselines in this paper. The output of ICRNN follows [17]:
$\begin{split}h_{t}&=g_{1}(U\textbf{x}_{t}+Wh_{t-1}+D_{2}\textbf{x}_{t-1})\\ y_{t}&=g_{2}(Vh_{t}+D_{1}h_{t-1}+D_{3}\textbf{x}_{t})\end{split}$
The output $y_{t}$ is convex with respect to the input x if non-decreasing and convex activation functions are used for $g$, and $[D_{3},D_{2},D_{1},V,W,U]$ are chosen to be non-negative weights, where x denotes $\big{[}\begin{smallmatrix}x\\ -x\end{smallmatrix}\big{]}$.
II-D Lipschitz RNN
@@
By adhering to the definition of Lipschitz continuity for a function $f$, where an L-Lipschitz continuous $f$ ensures that any minor perturbation to the input results in an output change of at most $L$ times the magnitude of that perturbation. Therefore, constraining neural networks to maintain Lipschitz continuity significantly fortifies their resilience against input perturbations.
A pivotal work, namely the Lipschitz RNN (LRNN), stands as one of the baselines in this paper. The output of LRNN follows [19]:
$\begin{split}\dot{h}&=A_{\beta_{A},\gamma_{A}}h+tanh(W_{\beta_{W},\gamma_{W}}h+Ux+b)\\ y&=Dh\end{split}$
where $\beta_{A},\beta_{W}\in[0,1]$, $\gamma_{A},\gamma_{W}>0$ are tunable parameters, $M_{A},M_{W},D,U$ are trainable weights, $\dot{h}$ is the time derivative of $h$ (i.e., $h$ can be updated by the explicit (forward) Euler scheme or a two-stage explicit Runge-Kutta scheme), and $A_{\beta_{A},\gamma_{A}},W_{\beta_{W},\gamma_{W}}$ is computed as follows:
$\begin{split}A_{\beta_{A},\gamma_{A}}=&(1-\beta_{A})(M_{A}+M_{A}^{T})+\beta_{A}(M_{A}-M_{A}^{T})\\ &-\gamma_{A}I\\ W_{\beta_{W},\gamma_{W}}=&(1-\beta_{W})(M_{W}+M_{W}^{T})+\beta_{W}(M_{W}-M_{W}^{T})\\ &-\gamma_{W}I\end{split}$
III Input Convex Lipschitz Recurrent Neural Networks
@@
In this section, we develop a novel network architecture, termed ICLRNN, that possesses both Lipschitz continuity and input convexity properties. Specifically, the output of ICLRNN is defined as follows:
@@
where all $W_{i}$ are constrained to be non-negative with singular values small and bounded by $\mathcal{O}(1)$, and all $g_{i}$ are constrained to be convex, non-decreasing, and Lipschitz continuous.
Similar to LRNN and ICRNN, UniCORNN and Linear Recurrent Units involve high Floating Point Operations (FLOPs). This aspect motivated us to retain the standard RNN architecture while imposing input convexity and Lipschitz constraints on the weights, since many engineering applications can achieve a satisfactory modeling accuracy even using a relatively simple NN structure. However, how to improve the robustness and computational efficiency of these NN models remains an important question (i.e., at this stage, we consider simple RNN as the starting point).
The output of a neural network inherits input convexity and Lipschitz continuity if and only if every hidden state possesses these properties. Unlike [17] and [19], our ICLRNN adheres to this foundational principle by imposing constraints on the weights and activation functions instead of introducing supplementary variables to the original RNN architecture. This approach minimizes the demand of computational resources. We first introduce a proposition to demonstrate that imposing an input convex constraint subsequent to a Lipschitz constraint does not compromise the latter.
The proofs of the following Propositions and Theorems can be found in Appendix A.
Proposition 1.
@@
For ICLRNN, we first enforce the spectral constraints to ensure that the largest singular values of the weight matrix is at most one [18]. The spectral constraint is implemented in two steps, which is similar to [20]. Firstly, we use spectral normalization as proposed by [25], where the spectral norm is reduced to at most 1 by iteratively evaluating the largest singular value with the power iteration algorithm proposed by [26]. Secondly, we apply the Björck algorithm proposed by [27] to increase other singular values to at most one.
Next, we enforce the non-negative constraint by performing weight clipping to set all negative values in the weights to 0. The theoretical upper bound after the integration of the input convex constraint with the Lipschitz constraint is $\sqrt{r(W)}$ for one single time step as shown in Proposition 1. It is essential to note that the upper bound in Eq. (6) is not precisely tight [28], and thus the upper bound developed in Proposition 1 is not tight. Through experimentation on random large matrices and the experiments in Section IV, we have observed that the Lipschitz constant is normally bounded within the order of $\mathcal{O}(1)$.
Moreover, ICLRNN uses ReLU as the activation function for hidden states instead of the GroupSort function proposed in [18]. This decision aims to maintain convexity within the model architecture. Although GroupSort exhibits gradient norm preservation and higher expressive power [10], its non-convex nature contrasts with our primary objectives. Next, we will proceed to prove the convexity and Lipschitz continuity of ICLRNN.
III-A Convexity and Lipschitz Continuity of ICLRNN
Proposition 2.
@@ III-A Convexity and Lipschitz Continuity of ICLRNN
By combining Proposition 2 and Proposition 3, the following theorem can be readily derived without requiring additional proof.
Theorem 1.
@@
Next, we proceed by developing a Lyapunov-based MPC (LMPC) framework. Our focus lies in showcasing the transformation of a non-convex NN-based LMPC into a convex optimization problem by utilizing ICNNs.
III-B ICLRNN for a Finite-Horizon Convex MPC
III-B1 Class of systems
Specifically, we examine a category of systems that can be expressed through a particular set of ordinary differential equations (ODE), given by the following form:
@@ III-B1 Class of systems
where $u\in\mathbb{R}^{m}$ denotes the control action and $x\in\mathbb{R}^{n}$ denotes the state vector. The function $F:X\times U\to\mathbb{R}^{n}$ is continuously differentiable, where $X\subset\mathbb{R}^{n}$ and $U\subset\mathbb{R}^{m}$ are connected and compact subsets that enclose an open neighborhood around the origin respectively. In this context, we maintain the assumption throughout this task that $F(0,0)=0$, ensuring that the origin $(x,u)=(0,0)$ serves as an equilibrium point. As it might be impractical to obtain first-principles models for intricate real-world systems, our objective is to construct a neural network to model the nonlinear system specified by Eq. (2), embed it into MPC, and maintain computational efficiency and non-adversarial robustness of MPC simultaneously.
III-B2 LMPC formulation
The optimization problem of the LMPC scheme, which incorporates a neural network model as its predictive element, is expressed as follows [29]:

### v4

@@ Abstract
Computational efficiency and non-adversarial robustness are critical factors in process modeling and optimization for real-world engineering applications. Yet, conventional neural networks often fall short in addressing both simultaneously, or even separately. Drawing insights from natural physical systems and existing literature, it is known theoretically that an input convex architecture will enhance computational efficiency, while a Lipschitz-constrained architecture will bolster non-adversarial robustness. However, integrating both properties into one model is a nontrivial task, as enforcing one property may compromise the other one. Therefore, in this work, we develop a novel network architecture, termed Input Convex Lipschitz Recurrent Neural Networks, that inherits the strengths of both convexity and Lipschitz continuity. This model is explicitly designed for fast and robust optimization-based tasks, which outperforms existing recurrent units in terms of computational efficiency and non-adversarial robustness. Additionally, we have successfully implemented this model in various practical engineering applications, such as optimization of chemical processes and real-world solar irradiance prediction for Solar PV system planning at LHT Holdings in Singapore. Source code is available at https://github.com/killingbear999/ICLRNN.
1 Introduction
Neural networks provide a powerful tool to process modeling, an important engineering problem in various manufacturing industries such as chemical, pharmaceutical, and energy systems. This work draws inspiration from nature-inspired design to develop a fast and robust solution that addresses two crucial challenges: computational efficiency and non-adversarial robustness, for modeling of real-world nonlinear systems. Specifically, nerual network models have been developed to capture complex nonlinear dynamics of industrial manufacturing systems, and the models can be further used for process monitoring, predictive maintenance, optimization, and real-time control. However, while neural network-based optimization can optimize process performance such as energy consumption and process profits, based on the prediction of neural networks, computational efficiency remains a big challenge for its real-time implementation. For example, model predictive control (MPC), an advanced optimization-based control technique, has been widely used to manage a process while complying with a set of constraints. Over the past decades, neural network-based MPC (NN-MPC) has received increasing attention, where a neural network that models system dynamics is incorporated into the design of MPC optimization problems [1, 2, 3, 4, 5, 6, 7, 8]. However, general optimization problems (e.g., MPC) built on conventional neural networks often fall short of meeting the requirement of computational efficiency. Nonetheless, efforts have been made to mitigate these limitations. For example, the computational efficiency of NN-MPC is enhanced by implementing an input convex architecture in neural networks [9]. Additionally, non-adversarial robustness of NN-MPC is improved by developing Lipschitz-constrained neural networks (LNNs) [10].
@@ 1 Introduction
In real-world applications, computational efficiency plays a pivotal role due to the imperative need for real-time decision-making, especially in critical engineering processes such as chemical process operations [9]. We define computational efficiency as the speed at which computational tasks are completed within a given execution timeframe. An effective approach to achieve this efficiency is through the transformation from non-convex to convex structures. Convexity is a ubiquitous characteristic observed in various physical systems such as the potential energy function of a mass on a spring. Building upon nature-inspired design and extending its application to neural networks, Input Convex Neural Networks (ICNNs) is a class of neural network architectures intentionally crafted to maintain convexity in their output with respect to the input. Leveraging these models could be particularly advantageous in various engineering problems, notably in applications such as MPC, due to their inherent benefits especially in optimization. This concept was first introduced as Feedforward Neural Networks (FNN) [11], and subsequently extended to Recurrent Neural Networks (RNN) [12], and Long Short-Term Memory (LSTM) [9].
@@
Additionally, in real-world engineering applications, non-adversarial robustness plays a critical role due to the prevalent noise present in most real-world data, which significantly hampers neural network performance. Our definition of robustness aligns closely with the principles outlined by [10]. Given the inherent noise in real-world process data, our goal is to improve neural networks by effectively learning from noisy training data for comprehensive end-to-end applications. Leveraging LNNs offers a potential solution to the robustness challenge. This concept was first introduced as FNN [13], and subsequently extended to RNN [14], and Convolutional Neural Networks (CNN) [15]. However, at this stage, the development of neural networks that address the above two challenges simultaneously remains an open question. It is important to note that the integration of ICNNs and FNNs presents a non-trivial challenge due to their interplay in the way that enforcing one property in the design of neural networks may compromise the other property.
@@
Motivated by the above considerations, in this work, we introduce an Input Convex Lipschitz RNN (ICLRNN) that amalgamates the strengths of ICNNs and LNNs. This approach aims to achieve the concurrent resolution of computational efficiency and non-adversarial robustness for real-world engineering applications. Specifically, we validate the efficacy of ICLRNN across various engineering problems that include neural network-based optimization for chemical processes and time series forecasting in a real-world Solar Photovoltaic (PV) system in Singapore. Performance assessment hinges on modeling accuracy, computational efficiency (i.e., evaluated by runtime and floating-point operations), and non-adversarial robustness. Our objective is to achieve equilibrium across these three dimensions as they hold equal importance in real-world engineering scenarios.
@@
In summary, Section 2 offers an overview of modern high-performing recurrent units, ICNNs and LNNs. Section 3 presents the proposed ICLRNN, substantiating its theoretical attributes as input convex and Lipschitz continuous. Section 4 illustrates that ICLRNN surpasses state-of-the-art recurrent units in various engineering tasks, including optimization and control of chemical processes using MPC and real-time solar irradiance forecasting on Solar PV systems. Section 5 delves into the potential limitations of ICLRNN uncovered through experiments conducted on the Long Range Arena benchmark [16].
2 Background
2.1 Notations
@@ 2.2 High-performing RNNs
In recent years, several high-performing recurrent units have emerged, including UniCORNN [17], rank-coding based RNNs [18], and Linear Recurrent Units [19]. However, these variants primarily emphasize enhancing accuracy, especially for very long time-dependent sequences modeling, which often at the expense of model simplicity and computational efficiency. Moreover, in many engineering applications that we are currently working on with industry partners, accuracy alone does not dictate the suitability of a model. In fact, from an engineering point of view, simplicity and robustness are also important in addition to accuracy, as they ensure the model can be easily adapted and incorporated into existing industrial operational systems, and can be trusted in an uncertain environment with noise and disturbances. Therefore, the goal of this work is to develop a network architecture based on simple RNNs that strikes a balance between these factors while maintaining a satisfactory level of accuracy, rather than investigating the novel designs of RNNs for improved accuracy only.
2.3 Input Convex RNN
@@
The initiative behind ICNNs is to leverage the power of neural networks, specifically designed to maintain convexity within their decision boundaries [11]. ICNNs aim to combine the strengths of neural networks in modeling complex data with the advantages of convex optimization, which ensures the convergence to a global optimal solution. They are particularly attractive for control and optimization problems, where achieving globally optimal solutions is essential. By maintaining convexity, ICNNs help ensure that the optimization problems associated with neural networks (e.g., neural network-based optimization) remain tractable, thus addressing some of the challenges of non-convexity and the potential for suboptimal local solutions in traditional neural networks. A foundational work, known as Input Convex RNN (ICRNN) [12], serves as one of the baselines.
2.4 Lipschitz RNN
@@
By adhering to the definition of Lipschitz continuity for a function $f$, where an $L$-Lipschitz continuous $f$ ensures that any minor perturbation to the input results in an output change of at most $L$ times the magnitude of that perturbation. Therefore, constraining neural networks to maintain Lipschitz continuity significantly fortifies their resilience against input perturbations. A pivotal work, namely the Lipschitz RNN (LRNN) [14], stands as one of the baselines in this paper.
3 Input Convex Lipschitz Recurrent Neural Networks
3.1 Convexity and Lipschitz Continuity of ICLRNN
@@
We propose a novel network architecture, termed ICLRNN, that possesses both Lipschitz continuity and input convexity properties. Specifically, the output of ICLRNN is defined as follows:
@@
where all $W_{i}$ are constrained to be non-negative with singular values small and bounded by $\mathcal{O}(1)$, and all $g_{i}$ are constrained to be convex, non-decreasing, and Lipschitz continuous. Similar to LRNN and ICRNN, UniCORNN and Linear Recurrent Units involve high Floating Point Operations, and do not provide a theoretical guarantee on non-adversarial robustness and computational efficiency on convex neural-network-based optimization problem. These aspects motivated us to retain the standard RNN architecture while imposing input convexity and Lipschitz constraints on the weights, since many engineering applications can achieve satisfactory modeling accuracy even using relatively simple NN structures. However, how to improve the robustness and computational efficiency of these NN models remains an important question.
The output of a neural network inherits input convexity and Lipschitz continuity if and only if every hidden state possesses these properties. Unlike [12] and [14], our ICLRNN adheres to this foundational principle by imposing constraints on the weights and activation functions instead of introducing supplementary variables to the original RNN architecture. This approach minimizes the demand of computational resources. The following proposition demonstrates that following our proposed method of imposing an input convex constraint subsequent to a Lipschitz constraint does not compromise the latter, following the proposed design of ICLRNNs. The proofs of the following Propositions and Theorems can be found in Appendix A.
We first establish Proposition 1 to demonstrate that the imposition of an Input Convex constraint subsequent to a Lipschitz constraint does not compromise the Lipschtiz property to a significant extent, allowing both to harmoniously coexist.
Proposition 1
@@
For ICLRNN, we first enforce the spectral constraints to ensure that the largest singular values of the weight matrix is at most one [13]. The spectral constraint is implemented in two steps, which is similar to [15]. Firstly, we use spectral normalization as proposed by [20], where the spectral norm is reduced to at most 1 by iteratively evaluating the largest singular value with the power iteration algorithm, which is a method used to find the dominant eigenvector of a square matrix using an iterative numerical technique that repeatedly multiplies the matrix by a vector and normalizes the result until convergence. Secondly, we apply the Björck algorithm, which is an iterative algorithm for computing the best estimate of an orthogonal matrix (i.e., iteratively refining the initial estimate of the orthogonal matrix by minimizing the error between the estimate and the true orthogonal matrix, eventually converging to a stable solution), to increase other singular values to at most one.
Next, we enforce the non-negative constraint by performing weight clipping to set all negative values in the weights to 0. The theoretical upper bound after the integration of the input convex constraint with the Lipschitz constraint is $\sqrt{r(W)}$ for one single time step as shown in Proposition 1. It is essential to note that the upper bound (i.e., $\|A\|=\sigma_{max}(A)\leq\|A\|_{F}$) is not precisely tight [21], and thus the upper bound developed in Proposition 1 is not tight. Through experimentation on random large matrices and the experiments in Section 4, we have observed that the Lipschitz constant is normally bounded within the order of $\mathcal{O}(1)$.
Moreover, ICLRNN uses ReLU as the activation function for hidden states instead of the GroupSort function proposed in [13]. This decision aims to maintain convexity within the model architecture. Although GroupSort exhibits gradient norm preservation and higher expressive power [10], its non-convex nature contrasts with our primary objectives. Next, by combining Proposition 1, Proposition 2, and Proposition 3, Theorem 1 indicates the convexity and Lipschitz continuity of ICLRNN.
Proposition 2
@@
Theorem 1
@@
To demonstrate the benefits of ICLRNN in neural-network-based optimization tasks, we proceed by constructing an MPC framework, one of the most widely-used control methods in chemical industries that regulates the process states to the desired set point by solving optimization problems in real time. The goal is to show that a non-convex NN-based optimization problem can be converted into a convex optimization problem through the utilization of ICNNs.
3.2 ICLRNN for a finite-horizon convex MPC
3.2.1 Class of systems
We consider a category of systems that can be expressed through the ordinary differential equations (ODEs) of the following form:
@@ 3.2.1 Class of systems
where $u\in\mathbb{R}^{m}$ denotes the control action and $x\in\mathbb{R}^{n}$ denotes the state vector. The function $F:X\times U\to\mathbb{R}^{n}$ is continuously differentiable, where $X\subset\mathbb{R}^{n}$ and $U\subset\mathbb{R}^{m}$ are connected and compact subsets that enclose an open neighborhood around the origin respectively. In this context, we maintain the assumption throughout this task that $F(0,0)=0$, ensuring that the origin $(x,u)=(0,0)$ serves as an equilibrium point. As it might be impractical to obtain first-principles models for intricate real-world systems, our objective is to construct a neural network to model the nonlinear system specified by Eq. (2), embed it into optimization problems, and maintain computational efficiency and non-adversarial robustness simultaneously.
3.2.2 MPC formulation
The optimization problem of the Lyapunov-based MPC (LMPC) that incorporates a ICLRNN as its predictive model and uses a Lyapunov function as stability constraints, is expressed as follows: